# Block Vecchia Approximation for Scalable and Efficient Gaussian Process Computations

### v1

@@ Abstract
Gaussian Processes (GPs) are vital for modeling and predicting irregularly-spaced, large geospatial datasets. However, their computations often pose significant challenges in large-scale applications. One popular method to approximate GPs is the Vecchia approximation, which approximates the full likelihood via a series of conditional probabilities. The classical Vecchia approximation uses univariate conditional distributions, which leads to redundant evaluations and memory burdens. To address this challenge, our study introduces block Vecchia, which evaluates each multivariate conditional distribution of a block of observations, with blocks formed using the K-means algorithm. The proposed GPU framework for the block Vecchia uses varying batched linear algebra operations to compute multivariate conditional distributions concurrently, notably diminishing the frequent likelihood evaluations. Diving into the factor affecting the accuracy of the block Vecchia, the neighbor selection criterion is investigated, where we found that the random ordering markedly enhances the approximated quality as the block count becomes large. To verify the scalability and efficiency of the algorithm, we conduct a series of numerical studies and simulations, demonstrating their practical utility and effectiveness compared to the exact GP. Moreover, we tackle large-scale real datasets using the block Vecchia method, i.e., high-resolution 3D profile wind speed with a million points.
@@ 1 Introduction
Gaussian Processes (GPs) play a crucial role in spatial statistics applications, where they are employed for modeling and predicting geospatial data. However, a significant challenge arises when dealing with large datasets collected at irregularly-spaced locations where the computational complexity of the GP modeling and prediction increases cubically with the number of spatial locations. This poses a limitation for applications with large spatial datasets. Indeed, assume there are $n$ locations ${\bf s}_{1},\ldots,{\bf s}_{n}\in\mathbb{R}^{d}$, where $d$ is the dimension of locations. The GPs are specified by a mean function and a covariance function, $GP(\mu({\bf s}),C_{\bm{\theta}}(\cdot,\cdot))$, where $C_{\bm{\theta}}$ is usually assumed to have a parametric form and $\bm{\theta}\in\mathbb{R}^{p}$. Let $y_{i}:=y\left({\bf s}_{i}\right)\in\mathbb{R}$ represent single observation at location ${\bf s}_{i}$ and denote the observed data vector by $\boldsymbol{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$. The data $\boldsymbol{y}$ can be modeled using GPs as $\boldsymbol{y}\sim\mathcal{N}_{n}\left(\bm{\mu},\bm{\Sigma}_{\boldsymbol{\theta}}\right)$, where $\bm{\Sigma}_{\boldsymbol{\theta}}$ is a covariance matrix with $(i,j)$ entry determined by a given covariance function $C_{\boldsymbol{\theta}}\left({\bf s}_{i},{\bf s}_{j}\right)$. Without loss of generality, we assume that $\boldsymbol{y}$ has a mean of zero. Statistical inference about $\boldsymbol{\theta}$ often relies on the Gaussian log-likelihood function
@@
Various research projects have focused on the computational challenges of modeling and predicting using large-scale GPs. These efforts primarily explore two strategies: sparse approximation and low-rank approximation of the covariance matrix (Furrer et al.,, 2006; Kaufman et al.,, 2008; Bevilacqua et al.,, 2016). For instance, sparse approximation techniques such as covariance tapering have been widely studied. This method applies a tapering function that diminishes to zero with increasing distance between two points to the covariance function. The result transforms the original dense covariance matrix into a more manageable sparse format. In low-rank approximation, the full covariance matrix is approximated with two matrices of lower rank, which is achieved by $\bf A\approx UV^{\top}$, where $\bf U$ and $\bf V$ are matrices with fewer columns than $\bf A$. This reduces the rank of the matrix to the number of columns in $\bf U$ and $\bf V$, thus significantly reducing the size of the matrix operations involved (Huang and Sun,, 2018; Mondal et al.,, 2023). Additionally, modern hardware technologies capable of low-precision calculations, such as NVIDIA GPUs, have enhanced the optimization of the sparse covariance matrix. This improvement is achieved by assigning different precision levels to various sections of the dense covariance matrix, thus reducing computational complexity rather than completely omitting these sections (Abdulah et al.,, 2019, 2021; Cao et al.,, 2022). For low-rank approximations, various methods are employed to enable faster computations and reduced memory compared to the original dense matrix (Katzfuss and Cressie,, 2011; Huang and Sun,, 2018; Abdulah et al.,, 2018b; Mondal et al.,, 2022).
@@
Vecchia approximation is one of the earliest GP statistical approximation methods. It involves replacing the high-dimensional joint distribution of the GP with a product of univariate independent conditional distributions. A small set of observations is selected as the conditioning set (Vecchia,, 1988). This method involves approximating each univariate conditional distribution without using all the neighbors, thus allowing faster execution and less memory consumption. Vecchia approximation results in an approximated log-likelihood function with a computational complexity of $\mathcal{O}(nm^{3})$ instead of the regular $\mathcal{O}(n^{3})$ complexity. Here, $n$ denotes the number of spatial locations, and $m$ represents the number of neighbors considered in the conditional distributions ($m\ll n$). Another advantage of the Vecchia approximation is that it is well-suited for parallel computing since each univariate conditional term can be computed independently. Nonetheless, the scalability of the Vecchia approximation encounters two primary hurdles. Firstly, the decomposition in univariate conditional distributions may lead to the redundant computation of multiple conditional log-likelihoods, particularly when they involve common neighbors. This redundancy not only increases computational overhead but also decreases the efficiency of processing. Secondly, the approximation’s reliance on small matrix operations, which are more suitably executed on CPUs rather than Graphics Processing Units (GPUs), represents a significant challenge in establishing a GPU framework of Vecchia approximation. GPUs are designed to tackle problems with large computational demands, typically suited for tasks that benefit from massive parallelization, such as matrix-to-matrix multiplications. In contrast, the nature of tasks generated by the Vecchia approximation, i.e., numerous small, discrete operations, does not align well with the architectural strengths of GPUs. For further reviews in Vecchia approximation, see (Pan et al.,, 2024; Katzfuss et al.,, 2022; Zhang and Katzfuss,, 2022; Guinness,, 2021, 2018; Katzfuss and Guinness,, 2021; Zhang et al.,, 2021; Jimenez and Katzfuss,, 2023).
@@
In this work, we introduce a block version of Vecchia approximation using the batched GPU framework, i.e., block Vecchia algorithm, where the full likelihood is approximated by the product of a series of multivariate conditional probabilities represented by blocks and their neighbors, and the modern GPU architectures are used to accelerate the proposed algorithm. Our approach, facilitated by relying on the MAGMA library (Dong et al.,, 2016), employs a strategy of approximating the full likelihood by the product of smaller multivariate conditional distributions, each of which indicates the multivariate distribution of a block of observations conditioning on their selected neighbors. We refer to these clusters as blocks, which are created by employing the K-means algorithm to group nearby locations together. Classical Vecchia can be viewed as a special case when the block size is 1. The enhancement in our GPU framework lies in applying varying batched matrix operations, which execute the computationally light tasks, e.g., the Cholesky decomposition of the small covariance matrix, in parallel on a single GPU by creating multiple block threads. Further investigation establishes that the block Vecchia approximation algorithm significantly enhances the capability to manage larger problem sizes, outperforming the classic Vecchia algorithm’s scalability. Through comprehensive evaluations encompassing numerical studies, simulation experiments, and analyses of real datasets in the context of the Matérn covariance function considered in the GP modeling, key findings have been identified, underscoring the efficiency and effectiveness of the block Vecchia method:
@@ 2 Block Vecchia Framework
This section offers an in-depth description of our proposed framework for the block Vecchia algorithm. It commences by explaining the preprocessing step, which entails location clustering, block reordering, and nearest neighbor searching for each centroid of blocks. Subsequently, we delve into the memory requirements for the block Vecchia algorithm with a comprehensive description of the proposed implementation. Then, the prediction using the block Vecchia algorithm is introduced. Finally, we compare the expected memory usage and computational complexity of the block Vecchia algorithm in contrast to the classic Vecchia algorithm.
@@ 2.3 Block Vecchia Prediction
The output, $\boldsymbol{y}_{*}$ and $\{\mathbf{\Sigma}_{B^{*}_{i}}\}_{i=1}^{bc}$, helps the univariate (and multivariate) conditional simulation and the prediction interval. Considering the simplicity and computational efficiency, the univariate configuration is adopted. The variance $\boldsymbol{\sigma}=\left(\sigma_{1},\sigma_{2},\cdots,\sigma_{n*}\right)^{\top}$ is extracted from the diagnal element from the $\{\mathbf{\Sigma}_{B^{*}_{i}}\}_{i=1}^{bc}$ for the predicted values $\boldsymbol{y}_{*}$. Then the conditional simulations are conducted using the $\boldsymbol{y}_{*}$ and $\boldsymbol{\sigma}$. We calculate the sample mean and variance $(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\sigma}}^{2})$ and then the 95% confidence interval is $\left(\tilde{\mu}-z_{\alpha/2}\tilde{\sigma}_{i},\tilde{\mu}+z_{\alpha/2}\tilde{\sigma}_{i}\right)$ with $\alpha=0.05$.
@@ 2.4 GPU and Batched Linear Algebra
Batched operations involve processing multiple data points or operations simultaneously, thereby reducing the overhead associated with individual computations (Haidar et al.,, 2015). In the context of the block Vecchia approximation, which consists of abundant independent and computationally light conditional probabilities, batched operations can be employed to improve computational efficiency by processing those covariance matrices and vectors at the same time, e.g., lines 17 - 21 and 24 - 27 in the Algorithm 1 in S1. In other words, when implemented on a GPU, these matrices and vectors can be processed in parallel batches, rather than sequentially as on a CPU, which dramatically reduces the computation overheads. By utilizing GPU acceleration through batched operations, the computational burden of the block Vecchia approximation is significantly reduced, allowing for faster processing times, especially when dealing with large datasets. This not only improves the feasibility of applying such methods in practice but also opens the door to analyzing more complex models and larger datasets that would be computationally prohibitive on traditional CPU-based systems.
@@ 2.5 Computational and Memory Complexity
We examine the memory footprint and the arithmetic complexity of the block Vecchia implementation, comparing it to the classic Vecchia implementation. The memory footprint of the classic Vecchia algorithm is $\mathcal{O}(nm^{2})$ for many small symmetric covariance matrices $\boldsymbol{\Sigma}_{i},i=1,\ldots,n$ and $\mathcal{O}(nm)$ for the conditioning vectors $\boldsymbol{y}_{NN,B^{\zeta}_{i}},i=1,\ldots,n$ where $m$ and $n$ stands for the conditioning size and observations, respectively. For the block Vecchia algorithm, each block requires three covariance matrices $\boldsymbol{\Sigma}_{i}^{lk}$, $\boldsymbol{\Sigma}_{i}^{cross}$, $\boldsymbol{\Sigma}_{i}^{con}$, $i=1,\ldots,bc$ where the average block size is approximately $n/bc$, and two observation vectors $\boldsymbol{y}_{NN,B^{\zeta}_{i}}$ and $\boldsymbol{y}_{B_{i}}$, $i=1,\ldots,bc$, with memory complexities of $\sim{bc(n/bc)^{2}/2}$, $\sim{bc*m(n/bc)/2}$ $\sim{bc*m^{2}/2}$, $\sim{bc*m}$, and $\sim{bc*(n/bc)}$, respectively. Therefore, the memory for the classic Vecchia is approximately $\sim{nm^{2}/2+nm}$ and for the block Vecchia is $\sim{n^{2}/(2bc)+mn/2}$ ${+m*(bc)^{2}/2+m*bc+n}$. In terms of the arithmetic complexity, for the block Vecchia, the complexity primarily stems from the Cholesky factorization of the covariance matrix, $\sim{bc(n/bc)^{3}/3}$ and $\sim{bc(m)^{3}/3}$, and matrix-to-matrix multiplication $\sim{2*bc*m(n/bc)^{2}}$. In contrast, with the classic Vecchia approximation, the complexity for the Cholesky factorization operations is $\sim{n(m^{3}/3)}$.
Figure 4 (a) shows the memory footprint in gigabytes (GB) for various problem sizes when employing the classic Vecchia and block Vecchia algorithms, which exhibit notably lower memory requirements for block Vecchia. Figure  4 (b) demonstrates the floating-point operations (flops) in Gflops for various problem sizes when using both the block Vecchia and classic algorithms. The figure shows a significant reduction in the number of required flops for the Vecchia algorithm, underlining its efficiency, particularly if it maintains accuracy comparable to the classic Vecchia algorithm.
@@
Figure 4: Comparison of Arithmetic complexity: block Vecchia (BV) versus classic Vecchia (CV) algorithms. The format of the legend is Method-BlockCount-ConditioningSize, e.g., CV-20K-30 representing the classic Vecchia with block count 20K and conditioning size 30; BV-2K-90 representing the block Vecchia with block count 2K and conditioning size 90.

### v2

@@ Abstract
Gaussian Processes (GPs) are vital for modeling and predicting irregularly-spaced, large geospatial datasets. However, their computations often pose significant challenges in large-scale applications. One popular method to approximate GPs is the Vecchia approximation, which approximates the full likelihood via a series of conditional probabilities. The classical Vecchia approximation uses univariate conditional distributions, which leads to redundant evaluations and memory burdens. To address this challenge, our study introduces block Vecchia, which evaluates each multivariate conditional distribution of a block of observations, with blocks formed using the K-means algorithm. The proposed GPU framework for the block Vecchia uses varying batched linear algebra operations to compute multivariate conditional distributions concurrently, notably diminishing the frequent likelihood evaluations. Diving into the factor affecting the accuracy of the block Vecchia, the neighbor selection criterion is investigated, where we found that the random ordering markedly enhances the approximated quality as the block count becomes large. To verify the scalability and efficiency of the algorithm, we conduct a series of numerical studies and simulations, demonstrating their practical utility and effectiveness compared to the exact GP. Moreover, we tackle large-scale real datasets using the block. Vecchia method, i.e., high-resolution 3D profile wind speed with a million points.
(Code: https://github.com/paper-code1/BV-Gaussian)
@@ 1 Introduction
Gaussian Process (GP) is an essential tool in spatial statistics, widely used for modeling and predicting geospatial data. Nonetheless, there is a major challenge when handling large datasets from irregularly distributed locations. Indeed, assume there are $n$ locations ${\bf s}_{1},\ldots,{\bf s}_{n}\in\mathbb{R}^{d}$, where $d$ is the dimension of locations. The GPs are specified by a mean function and a covariance function, $GP(\mu({\bf s}),C_{\bm{\theta}}(\cdot,\cdot))$, where $C_{\bm{\theta}}$ is usually assumed to have a parametric form and $\bm{\theta}\in\mathbb{R}^{p}$. Let $y_{i}:=y\left({\bf s}_{i}\right)\in\mathbb{R}$ represent single observation at location ${\bf s}_{i}$ and denote the observed data vector by $\boldsymbol{y}=\left(y_{1},\ldots,y_{n}\right)^{\top}$. The data $\boldsymbol{y}$ can be modeled using GPs as $\boldsymbol{y}\sim\mathcal{N}_{n}\left(\bm{\mu},\bm{\Sigma}_{\boldsymbol{\theta}}\right)$, where $\bm{\Sigma}_{\boldsymbol{\theta}}$ is a covariance matrix with $(i,j)$ entry determined by a given covariance function $C_{\boldsymbol{\theta}}\left({\bf s}_{i},{\bf s}_{j}\right)$. Without loss of generality, we assume that $\boldsymbol{y}$ has a mean of zero. Statistical inference about $\boldsymbol{\theta}$ often relies on the Gaussian log-likelihood function
@@
Various studies have focused on the computational and memory challenges of modeling and predicting using large-scale GPs. These efforts primarily explore two strategies: sparse approximation and low-rank approximation of the covariance matrix (Furrer et al.,, 2006; Kaufman et al.,, 2008; Bevilacqua et al.,, 2016). For instance, sparse approximation techniques such as covariance tapering have been widely studied. This method applies a tapering function that diminishes to zero with increasing distance between two points to the covariance function. This function transforms the original dense covariance matrix into a sparse format, thus reducing the computations and memory burden. In low-rank approximation, the full covariance matrix is approximated with two matrices of lower rank, which is achieved by $\bf A\approx UV^{\top}$, where $\bf U$ and $\bf V$ are matrices with fewer columns than $\bf A$. This reduces the rank of the matrix to the number of columns in $\bf U$ and $\bf V$, thus significantly reducing the size of the matrix operations involved (Huang and Sun,, 2018; Mondal et al.,, 2023). Additionally, modern hardware technologies capable of low-precision calculations, such as NVIDIA GPUs, have enhanced the optimization of the sparse covariance matrix. This improvement is achieved by assigning different precision levels to various sections of the dense covariance matrix, thus reducing computational complexity rather than completely omitting these sections (Abdulah et al.,, 2019, 2021; Cao et al.,, 2022). For low-rank approximations, various methods are employed to enable faster computations and reduced memory compared to the original dense matrix (Katzfuss and Cressie,, 2011; Huang and Sun,, 2018; Abdulah et al.,, 2018b; Mondal et al.,, 2022).
@@
The Vecchia approximation is among the earliest statistical methods for approximating GPs. It approximates the joint distribution of a GP by decomposing it into a product of independent univariate conditional distributions(Vecchia,, 1988). This approach reduces computational demands and memory burden by using only a limited number of neighboring points in each univariate conditional distribution, improving speed and lowering memory requirements. The approximated log-likelihood has a computational complexity of $\mathcal{O}(nm^{3})$ and memory complexity of $\mathcal{O}(nm^{2})$, compared to the exact $\mathcal{O}(n^{3})$ and $\mathcal{O}(n^{2})$. Here, $n$ is the total number of spatial locations, and $m$ is the number of neighbors included in the conditioning set ($m\ll n$). Nonetheless, the scalability of the Vecchia approximation encounters two primary hurdles. Firstly, the decomposition in univariate conditional distributions may lead to the redundant computation of multiple conditional log-likelihoods, particularly when they involve common neighbors. This redundancy not only increases computational overhead but also decreases the efficiency of processing. Secondly, the main computations in the Vecchia approximation are small matrix operations, which are better suited for execution on CPUs than Graphics Processing Units (GPUs). More importantly, the performance of GPUs cannot be fully utilized for conducting the small matrix operations(Pan et al.,, 2024). GPUs are designed for handling computationally intensive tasks that benefit from extensive parallelization, such as matrix-to-matrix multiplications. However, the tasks arising from the Vecchia approximation which is characterized by numerous small and discrete operations are not inherently compatible with the GPU architecture’s strengths. For a detailed review of the Vecchia approximation, refer to(Pan et al.,, 2024; Katzfuss et al.,, 2022; Zhang and Katzfuss,, 2022; Guinness,, 2021, 2018; Katzfuss and Guinness,, 2021; Zhang et al.,, 2021; Jimenez and Katzfuss,, 2023).
@@
In this work, we introduce a block version of Vecchia approximation using the batched GPU framework, i.e., block Vecchia algorithm, where the full likelihood is approximated by the product of a series of multivariate conditional probabilities represented by blocks and their neighbors, and modern GPU architectures are used to accelerate the proposed algorithm. These blocks are created by the K-means algorithm and the classic Vecchia can be viewed as a special case when each block size is 1. Technically, built upon the MAGMA library (Dong et al.,, 2016), the enhancement in our GPU framework lies in applying varying batched matrix operations, which execute the computationally light tasks, e.g., the Cholesky decomposition of the small covariance matrix, in parallel on a single GPU by creating multiple block threads. The batched operations accelerate the algorithm and the consideration of multivariate conditional probability creates computationally intensive tasks, which makes the compute node in GPU more efficient compared to univariate conditional cases in the classic Vecchia. Additionally, further investigation establishes that the block Vecchia approximation algorithm significantly enhances the capability to manage larger problem sizes, outperforming the classic Vecchia algorithm’s scalability. In the end, through comprehensive evaluations encompassing numerical studies, simulation experiments, and analyses of real datasets in the context of the Matérn covariance function considered in the GP modeling, key findings have been identified, underscoring the efficiency and effectiveness of the block Vecchia method:
@@ 2 Block Vecchia Framework
This section provides an overview of the proposed framework for the block Vecchia algorithm. It begins with an explanation of the preprocessing steps, including location clustering, block reordering, and nearest neighbor selection for each block centroid. Following this, the memory requirements for the block Vecchia algorithm are discussed in detail, alongside a thorough description of the proposed implementation. Next, the prediction process using the block Vecchia algorithm is outlined. Finally, a comparison is presented, highlighting the expected memory usage and computational complexity of the block Vecchia algorithm relative to the classic Vecchia algorithm.
@@ 2.3 Block Vecchia Prediction
The output, $\boldsymbol{y}_{*}$ and $\{\mathbf{\Sigma}_{B^{*}_{i}}\}_{i=1}^{bc}$, helps the univariate (and multivariate) conditional simulation and the prediction interval. Considering the simplicity and computational efficiency, the univariate configuration is adopted. The variance $\boldsymbol{\sigma}=\left(\sigma_{1},\sigma_{2},\cdots,\sigma_{n*}\right)^{\top}$ is extracted from the diagnal element from the $\{\mathbf{\Sigma}_{B^{*}_{i}}\}_{i=1}^{bc}$ for the predicted values $\boldsymbol{y}_{*}$. Then the conditional simulations are conducted using the $\boldsymbol{y}_{*}$ and $\boldsymbol{\sigma}$. We calculate the sample mean and variance $(\tilde{\boldsymbol{\mu}},\tilde{\boldsymbol{\sigma}}^{2})$ and then the 95% confidence interval is $\left(\tilde{\mu}-z_{\alpha/2}\tilde{\sigma}_{i},\tilde{\mu}+z_{\alpha/2}\tilde{\sigma}_{i}\right)$ with $\alpha=0.05$. For general applications, we propose the following scheme. Firstly, it is recommended to use a small block count and conditioning size to estimate the parameters. Then, these estimated parameters serve as initial values for accurate settings with larger block counts and conditioning sizes, continuing until a clear convergence trend emerges. This approach eliminates the need to determine the optimal block count and conditioning size explicitly, as real-world applications often vary in problem size and required precision. Employing this automated estimation method provides a more adaptable and practical solution.
@@ 2.4 GPU and Batched Linear Algebra
Batched operations involve processing multiple data points or operations simultaneously, thereby reducing the overhead associated with individual computations (Haidar et al.,, 2015). In the context of the block Vecchia approximation, which consists of abundant independent and computationally light conditional probabilities, batched operations can be employed to improve computational efficiency by processing those covariance matrices and vectors at the same time, e.g., lines 17 - 21 and 24 - 27 in the Algorithm 1 in S1. In other words, when implemented on a GPU, these matrices and vectors can be processed in parallel batches rather than sequentially as on a CPU, which dramatically reduces the computation overheads. By utilizing GPU acceleration through batched operations, the computational burden of the block Vecchia approximation is significantly reduced, allowing for faster processing times, especially when dealing with large datasets. This not only improves the feasibility of applying such methods in practice but also opens the door to analyzing more complex models and larger datasets that would be computationally prohibitive on traditional CPU-based systems.
@@ 2.5 Computational and Memory Complexity
We analyze the memory usage and computational complexity of the block Vecchia implementation, comparing it against the traditional Vecchia approach. The memory footprint of the classic Vecchia algorithm is $\mathcal{O}(nm^{2})$ for many small symmetric covariance matrices $\boldsymbol{\Sigma}_{i},i=1,\ldots,n$ and $\mathcal{O}(nm)$ for the conditioning vectors $\boldsymbol{y}_{NN,B^{\zeta}_{i}},i=1,\ldots,n$ where $m$ and $n$ stands for the conditioning size and observations, respectively. For the block Vecchia algorithm, each block requires three covariance matrices $\boldsymbol{\Sigma}_{i}^{lk}$, $\boldsymbol{\Sigma}_{i}^{cross}$, $\boldsymbol{\Sigma}_{i}^{con}$, $i=1,\ldots,bc$ where the average block size is approximately $n/bc$, and two observation vectors $\boldsymbol{y}_{NN,B^{\zeta}_{i}}$ and $\boldsymbol{y}_{B_{i}}$, $i=1,\ldots,bc$, with memory complexities of $\sim{bc(n/bc)^{2}/2}$, $\sim{bc*m(n/bc)/2}$ $\sim{bc*m^{2}/2}$, $\sim{bc*m}$, and $\sim{bc*(n/bc)}$, respectively. Therefore, the memory for the classic Vecchia is approximately $\sim{nm^{2}/2+nm}$ and for the block Vecchia is $\sim{n^{2}/(2bc)+mn/2}$ ${+bc*m^{2}/2+m*bc+n}$. In terms of the arithmetic complexity, for the block Vecchia, the complexity primarily stems from the Cholesky factorization of the covariance matrix, $\sim{bc(n/bc)^{3}/3}$ and $\sim{bc(m)^{3}/3}$, and matrix-to-matrix multiplication $\sim{2*bc*m(n/bc)^{2}}$. In contrast, with the classic Vecchia approximation, the complexity for the Cholesky factorization operations is $\sim{n(m^{3}/3)}$.
In Figure 4, to have a fair comparison between classic Vecchia and block Vecchia, we set the approximate block size 100, i.e., $n/bc\approx 100$ and nearest neighbor 3X larger than the classic Vecchia. Figure 4 (a) shows the memory footprint in gigabytes (GB) for increasing problem sizes when employing the classic Vecchia and block Vecchia algorithms. The block Vecchia notably lowers the memory burden. Figure  4 (b) demonstrates the floating-point operations (flops) in Gflops when using the block Vecchia and classic algorithms. The figure shows a comparable computation in the number of required flops for the Vecchia algorithm. The observations underscore that the block Vecchia creates computationally intensive tasks compared to the classic Vecchia and then it inspires us to utilize the modern GPU architecture to accelerate the computations.
@@
Figure 4: Comparison of Arithmetic complexity: block Vecchia (BV) versus classic Vecchia (CV) algorithms. The format of the legend is Method-ConditioningSize, e.g., CV-20K-30 represents the classic Vecchia with conditioning size 30; BV-2K-90 represents the block Vecchia with block count 2,000 and conditioning size 90.